OpenStudy (daddy_longlegs):

Please help Use the functions m(x) = 5x + 4 and n(x) = 6x − 9 to complete the function operations listed below. Part A: Find (m + n)(x). Show your work. Part B: Find (m ⋅ n)(x). Show your work. Part C: Find m[n(x)]. Show your work.

OpenStudy (pottersheep):

Another way to look at Part A (m+n)(x) is m(x)+n(x). So (5x+4) + (6x-9).

OpenStudy (pottersheep):

For Part B, which is (m⋅n)(x), another way to look at it is: m(x)⋅n(x) so, (5x+4)(6x-9)

OpenStudy (pottersheep):

For part c, which is m[n(x)] so m(x) = 5x + 4 now you have to sub in the function n(x) whereever there is an "x" in m(x). m(x) = 5x + 4 m(n(x)) = 5(n(x))+4 = 5(6x-9) + 4

OpenStudy (daddy_longlegs):

Thank you!

OpenStudy (pottersheep):

No problem :)